Cost Information and FIFO Gunnison Company had the following equivalent units schedule and cost information for its Sewing Department for the month of December: Direct Materials Conversion Costs Units started and completed 40,000 40,000 Add: Units in beginning work in process × Percentage complete: 5,000 × 0% direct materials — 5,000 × 50% conversion Costs 2,500 Add: Units in ending work in process × Percentage complete: 12,000 × 100% direct materials 12,000 — 12,000 × 35% conversion Costs — 4,200 Equivalent units of output 52,000 46,700 Costs: Work in process, December 1: Direct Material $65,000 Conversion Costs 10,000 Total work in process $75,000 Current costs: Direct Material $728,000 Conversion Costs 140,100 Total current costs $868,100

What if you were asked for the unit cost from the month of November? Calculate November's unit cost.
Answer #1

Ans

Cost per equivalent unit for November Month

17

1

Materials

14

=Current material cost/Equivalent units of Direct materials=728000/52000

Conversion

3

= Current Conversion cost/Equivalent units of conversion =140100/46700

Cost per equivalent unit for November Month

17

=14+3

Hope this helped ! Let me know in case of any queries.